OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points (5,2) and (-5,2).

jimthompson5910 (jim_thompson5910):

Notice how the y coordinates are the same. What do you think that means for the slope?

OpenStudy (anonymous):

No slope.

jimthompson5910 (jim_thompson5910):

close, but a slope does actually exist

OpenStudy (anonymous):

Ah, ok.

jimthompson5910 (jim_thompson5910):

what else do you think the slope is

OpenStudy (anonymous):

I would of originally said undefined slope.

jimthompson5910 (jim_thompson5910):

that's if the line was vertical and the x coordinates were equal

jimthompson5910 (jim_thompson5910):

but the slope isn't undefined in this case

OpenStudy (anonymous):

Ok. I'm sorry I just have trouble understanding all this type of math.

jimthompson5910 (jim_thompson5910):

that's ok, use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

There is a difference: \[\frac{0}{Anything( \ne 0)} \implies 0 (Defined)\] \[\frac{Anything}{0} \implies Undefined\]

OpenStudy (anonymous):

Oh fun.

OpenStudy (anonymous):

Firstly you should learn what are x and y coordinates in \((5,2)\).. Can you tell me what is x coordinate and what is y coordinate here ??

OpenStudy (anonymous):

So it would be: \[m=2-(-5)\]---------\[2-5\]

OpenStudy (anonymous):

Theslope is zero

OpenStudy (anonymous):

X = 5. Y = 2.

OpenStudy (anonymous):

Yes great.. Similarly what is x and y for other point given to you ??

OpenStudy (anonymous):

X = -5 and Y = 2.

OpenStudy (anonymous):

Yes.. Now subtract y coordinates of both the points..

OpenStudy (anonymous):

\[5--5=10\]

OpenStudy (anonymous):

But you said these are values of X ???

OpenStudy (anonymous):

I am saying to subtract Y coordinates..

OpenStudy (anonymous):

Oh sorry lol. 2 - 2 = 0

OpenStudy (anonymous):

Yes.. This will give you some hint about the answer but we have not done yet.. Now subtract the X coordinates..

OpenStudy (anonymous):

You will subtract in the same order as you did for Y coordinates..

OpenStudy (anonymous):

5 - -5=10

OpenStudy (anonymous):

Yes... Now divide both the answer you got..

OpenStudy (anonymous):

First answer you got is 0.. Secondly you got 10.. Divide them..

OpenStudy (anonymous):

10/0 = 0

OpenStudy (anonymous):

How ?? First answer you got was 0..

OpenStudy (anonymous):

First Answer divided by Second Answer..

OpenStudy (anonymous):

You can't divide by 0.

OpenStudy (anonymous):

But you are dividing this not I am..

OpenStudy (anonymous):

See if your first answer was 5 and Second Answer was 10 Then: First Answer divided by Second Answer means: 5/10 Not 10/5..

OpenStudy (anonymous):

Now tell me your first answer is 0 and second answer is 10 So what is First Answer divided by Second Answer ??

OpenStudy (anonymous):

It is a rule that you can not divide anything by 0 because 0 is basically nothing. And dividing 10 by 0 would be 0. Since 10 * 0 = 0. If you divide 0 by 10 then the answer is not possible. You never said anything either about dividing by 5, you just said to divide the answers, which was 0 and 10.

OpenStudy (anonymous):

Yes if you have 0 and 10 Then why are you dividing like 10/0 ?? Can't you divide like 0/10 ???

OpenStudy (anonymous):

I said to divide the answer in particular order..

OpenStudy (anonymous):

I did it both ways, one way it is not possible the other is 0. Check on the calculator if you don't believe me it will say the same thing. Like I just checked.

OpenStudy (anonymous):

Okay.. Just replace the question mark with the respective values: \[\frac{First \; Answer}{Second \;Answer} = \frac{?}{?}\] First answer is 0 and second answer is 10.. Plug in the values there..

OpenStudy (anonymous):

0 over 10.

OpenStudy (anonymous):

Yes.. Can you tell me what is the value of this ??

OpenStudy (anonymous):

0/10 = ??

OpenStudy (anonymous):

This I don't get cause 10*0=0 and 0*10=0

OpenStudy (anonymous):

Remember: 0 by anything is 0.. But anything must not be 0.. 0/10 = 0... This point you are getting ??

OpenStudy (anonymous):

0/1, 0/2, 0/3 etc all are 0.. But: 0/0 is not 0. Rather it is undefined..

OpenStudy (anonymous):

So your saying it is 10?

OpenStudy (anonymous):

Can you tell me where I have written 0/10 = 10 ??

OpenStudy (anonymous):

I think your eyes are somewhat weak : \[\Huge \color{green}{\frac{0}{Anything(except \; 0)} = 0}\]

OpenStudy (anonymous):

So now tell me what is the value of: \[\Huge \color{blue}{\frac{0}{10} = ??}\]

OpenStudy (anonymous):

With all due respect you saying my eyes are weak is rude. I have 20/20 vision thank you very much. And my problem is that your not explaining it to where I understand and your wanting to argue about being able to divide 0 by 10 when you can not.

OpenStudy (alexwee123):

@MathFrustration it is possible 0 is divisible 10

OpenStudy (alexwee123):

that * 0

OpenStudy (alexwee123):

0 divisible by 10*

OpenStudy (anonymous):

All calculator's said it is not possible.

OpenStudy (alexwee123):

 what do you mean? o.0